Grade 11Physical Chemistry

1.5 × 10^22 atomd of element weigh 0.9g . The atomic mass of the element (in amu)

Aryan
8 Years agoGrade 11

1 Answer

Surya
8 Years ago
Mol=N/Na=M/Mw=V/V0=1.5*10^22÷6.02*10^23=0.0250.025=0.9÷MwMw=0.9/0.25=900/25=180/5=36gSo the atomic mass of the given element is 36 g.